$SPAD/src/input explot3d.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Input for page ExPlot3DParametricCurve
)clear all
draw(curve(cos(t),sin(t),t),t=0..6)
draw(curve(t,t**2,t**3),t=-3..3)
-- Input for page ExPlot3DParametricSurface
)clear all
draw(surface(5*sin(u)*cos(v),4*sin(u)*sin(v),3*cos(u)),u=0..%pi,v=0..2*%pi)
draw(surface(u*cos(v),u*sin(v),u),u=0..4,v=0..2*%pi)
-- Input for page ExPlot3DFunctions
)clear all
draw(cos(x*y),x = -3..3,y = -3..3)

———-

References^{[1] nothing}

4